Direct Labor Cost Budget

MatchPoint Racket Company manufactures two types of tennis rackets, the Junior and Pro Striker models. The production budget for March for the two rackets is as follows:

Junior Pro Striker
Production budget 5,800 units 22,300 units

Both rackets are produced in two departments, Forming and Assembly. The direct labor hours required for each racket are estimated as follows:

Forming Department Assembly Department
Junior 0.25 hour per unit 0.5 hour per unit
Pro Striker 0.3 hour per unit 0.65 hour per unit

The direct labor rate for each department is as follows:

Forming Department $20 per hour
Assembly Department $10 per hour

Prepare the direct labor cost budget for March.

MatchPoint Racket Company
Direct Labor Cost Budget
For the Month Ending March 31
Forming
Department
Assembly
Department
Hours required for production:
Junior
Pro Striker
Total hours required
Hourly rate x$ x$
Total direct labor cost $ $
Answer #1
MatchPoint Racket Company
Direct Labor Cost Budget
For the Month Ending March 31
Forming Assembly
Department Department
Hours required for production:
Junior ((5800 units) = ( 0.25 *5800) = ( 0.5 *5800)
= 1450 = 2900
Pro Striker (22300 units) = ( 0.3 *22300) = ( 0.65 *22300)
= 6690 = 14495
Total hours required 8140 17395
Hourly rate $ 20 $ 10
Total direct labor cost $ 162800 $ 173950